I'm having a difficult time figuring out how to set up my equations in order to solve using a system of equations. The problem:
In winning the Western Conference finals, the San Antonia Spurs once scored 76 of their points on a combination of 36 two- and three- pointers. How many shots of each type did they make?
I initially set it up as:
x + y = 76,
3x + 2y = 36,
but that clearly didn't work out. I appreciate the help. Thank you!

Found 2 solutions by Boreal, rothauserc:
Answer by Boreal(15235)

x=2 point shots
y=3 point shots
x+y=36
2x+3y=76
-2x-2y=-72
y=4
x=32
That is 64 points from 2 point shots and 12 points from 3 point shots

Answer by rothauserc(4718)

3x + 2y = 76, where x is number of 3 point shots and y is number of 2 point shots
x + y = 36
solve for x and substitute in first equation
x = 36 - y
3(36 - y) + 2y = 76
108 -3y + 2y = 76
-y = -32
y = 32
x + 32 = 36
x = 4, therefore
there were 32 2-point shots and 4 3-point shots
************************************************
check the answer
3(4) + 2(32) = 76
12 + 64 = 76
76 = 76
our answer checks